Recall that the Natural Resources Defense Council found that 25% of bottled-water brands fill their bottles with just tap water. In a random sample of 70 bottled-water brands, let x be the number that contain tap water.

Find the approximate probability that 25 or more of the 70 sampled bottled-water brands will contain tap water.

The answer is .0192 but i am unsure how to solve

mean=17.5

std=3.62

zscore=2.07
Answer #1

Solution:

We are given

n = 70

p = 0.25

q = 1 – p = 0.75

Mean = np = 70*0.25 = 17.5

Standard deviation = SD = sqrt(npq) = sqrt(70*0.25*0.75) = 3.622844

WE have to find P(X>25)

Z = (X – mean)/SD = (25 - 17.5)/ 3.622844 = 2.070197

P(Z<2.070197) = P(X<25) = 0.980783

(by using z-table)

P(X>25) = 1 - P(X<25) = 0.019217

Answer: 0.0192